### v1

@@ 5.6 Can Generative AI Replace IF Staining Lab Work?
At current performance, generative AI cannot directly replace IF staining wet lab work, but it does show significant potential, particularly if issues surrounding generalisability can be addressed. In this study, we produced synthetic images of high visual quality that achieved comparable performance in predicting MoA in single-cell states. Notably, a visual assessment from an experienced lab scientist (eyeballing) found that the IF images generated by the Pix2Pix model were indistinguishable from real ones.
@@
The limited generalisability of these models does not make them entirely obsolete. Firstly, the limitation can potentially be addressed by scaling up the model; Large generative AI models, such as StableDiffusion, that have been trained on extensive, existing datasets have demonstrated generalisability to conditions the models had never seen before. It is likely this study does not approach the necessary size and scale required to produce such results. Secondly, synthetic IF images still hold significant potential in specific applications, such as longitudinal analysis and time-elapsed imaging, allowing for real-time, live imaging analysis.

### v2

@@ 5.5 Can Generative AI Replace IF Staining Lab Work?
At current performance, generative AI cannot directly replace IF staining wet lab work, but it does show significant potential, particularly if issues surrounding generalisability can be addressed. In this study, we produced synthetic images of high visual quality that achieved comparable performance in predicting MoA in single-cell states. Notably, a visual assessment from two experienced lab scientists (eyeballing) found that the IF images generated by the Pix2Pix model were indistinguishable from real ones.
@@
The lack of generalisability also introduces another issue: the potential learning curve for biologists to obtain synthesis results. However, we argue that this learning curve correlates with the generalisability of the synthesis models. Running inferences is straightforward if a model is already trained. For example, image synthesis tools like Stable Diffusion [1] and Midjourney [18] are easily used by individuals without a computer science background because these models are highly generalisable and can directly generate synthetic images or responses.
Currently, we acknowledge that IF synthesis models lack this level of generalisability. For unseen cell lines, biologists would need to train their own models, which is more challenging. This process involves curating a dataset, acquiring training resources, and evaluating the models. Therefore, the generalisability of the models significantly impacts the practicality of virtual IF staining.
The limited generalisability of these models does not make them entirely obsolete. Firstly, synthetic IF images still hold significant potential in specific applications, such as longitudinal analysis and time-elapsed imaging, allowing for real-time, live imaging analysis. Secondly, the limitation can potentially be addressed by scaling up the model; Large generative AI models, such as StableDiffusion [1], that have been trained on extensive, existing datasets have demonstrated generalisability to conditions the models had never seen before. It is likely this study does not approach the necessary size and scale required to produce such results.
@@
Last but not least, a significant concern regarding the use of generative AI in clinical settings is the ethical implications of using AI-generated images. The trustworthiness of these images can undermine the credibility of research papers that rely on synthesized data, since generated images are not "ground truths." The misuse of generative AI, including image manipulation and the creation of fake data, is another major concern.
5.6 General Generative AI for Biology: How Far Are We?
In the previous discussion section, we examined how the lack of generalisability affects the utility of generative AI in IF image generation. One forward solution for increasing generalisability is scaling up the generative AI models with larger parameters and extensive datasets.
However, several challenges need to be overcome before scaling up. First is the lack of high-quality datasets. High-quality datasets covering a wide range of cell types and conditions are often limited. For example, StableDiffusion was initially trained on 2.3 billion images [23]. In comparison, the entire cell painting repository [13] contains fewer than 50 million images, where only 220k images were used in our paper. Acquiring a dataset comparable to those used in natural image synthesis models is impractical for a single lab due to the immense cost and effort required. However, this does not mean the task is impossible. Collaborative efforts among multiple institutions can pool resources and data, making it feasible to build comprehensive datasets.
Second is the issue of limited resources. Training generative AI models typically requires significant computational resources, including high-performance GPUs and large memory capacities, which can be a barrier for many labs. The financial and infrastructural demands for such resources are substantial, making it challenging for smaller institutions to participate in this advanced research.